nancy made 76 and 58 on her first two algebra test and she has one more test in hte class. if she needs at least a 77 average to pass the course, what does she need to score on her third test to pass the course?
Answer by mananth(16946)

let her need x
Total will be (76+58+x )
Average = 77 of 3 tests
(76+58+x )/3 = 77
134+x=231
x=231-134
x=97
she needs 97